OpenStudy (anonymous):

x + y = 4 2x + 3y = 0

OpenStudy (greencat):

You need to subtract the equations.

OpenStudy (greencat):

But, before you do that you will have to multiply one equation.

OpenStudy (greencat):

Do you want to cancel out x or y?

OpenStudy (fibonaccichick666):

or solve one for each variable and substitute or create a coefficient matrix. Please decide which you prefer

OpenStudy (greencat):

Yes, you couls also use substitution; it's easier.

OpenStudy (greencat):

use 1st equation: x=4-y

OpenStudy (greencat):

plug it into 2nd equation: 2x+3y=0

OpenStudy (greencat):

2(4-y)+3y=0

OpenStudy (fibonaccichick666):

@GreenCat please allow @loveeaj22 to do the problem

OpenStudy (greencat):

yes sir

OpenStudy (greencat):

the next part is simple. give it a try.

OpenStudy (greencat):

yes ma'am*

OpenStudy (anonymous):

so i just plug in numbers for y and 3

OpenStudy (anonymous):

@GreenCat

OpenStudy (fibonaccichick666):

No, you do not have any numbers to plug in. You need to solve for y

OpenStudy (greencat):

2(4-y)+3y=0 multiply the 2 into the parenthesis.

OpenStudy (anonymous):

2x4=8-y ?

OpenStudy (fibonaccichick666):

please show us how you arrived there

OpenStudy (anonymous):

no wait i have the answer x=12 and y=-8 here ill check it for you (2x12)+ (3x-8)=0 12+-8=4

OpenStudy (anonymous):

i multiplied 2 on both sides and added it up then i pluged it into the first equation to check it @FibonacciChick666

OpenStudy (fibonaccichick666):

uhm well your check is correct, it works, but I still do not understand how you arrived at that point.

OpenStudy (greencat):

Why did you multiply 2 on both sides?

OpenStudy (anonymous):

because i thought that was what @GreenCat said but i guess i was wrong sorry !

OpenStudy (greencat):

Oops no! 2(4-y)+3y=0 ^ This 2 should be multiplied with 4-y